what is a2 + b2 = c2 if a = 12 and b = 12
Found 2 solutions by linh117, edjones:
Answer by linh117(31)

Pythagorean theorem ?
12^2 + 12^2 = c^2
144 + 144 = c^2
288= c^2
squareroot288=c
c=12root2

Answer by edjones(8007)

a^2 + b^2 = c^2
12^2+12^2=c^2
c^2=144*2
c=sqrt(144*2)
=12sqrt(2)
.
Ed